Solve this quadratic equation using the quadratic formula.
2x2 - 2x = 1

Answer by eperette(173)

CONCEPT YOU SHOULD KNOW:
QUADRATIC FORMULA (-b +- sqr(b^2-4ac))/2a
ax^2 + bx + c =0
***************
2x^2 - 2x =1
2x^2 -2x -1 =0
a= 2
b= -2
c= -1
(--2 +- sqr((-2)^2-4(2)(-1)))/ (2(2))
(2 +- sqr(4+8))/ 4
(2 +- sqr(12))/4
(2 +- sqr(4*3))/4
(2 +- 2sqr(3))/4
(1 +- 1sqr(3))/2
(1 + sqr (3))/2 or (1 - sqr(3))/2